OpenStudy (anonymous):

if 2(x-5)=-11 then x =

OpenStudy (anonymous):

Start by distributing the 2 to each term in the parenthesis.

OpenStudy (bobbyleary):

2(x-5)=-11 2x - 10 = -11 2x=-1 x=-1/2

OpenStudy (anonymous):

thk u

OpenStudy (bobbyleary):

No problem.